OpenStudy (anonymous):

If you write the following expression (x^−3*x^2/x^3*x^−2)^2 as a single power of x then the exponent is

OpenStudy (anonymous):

@zepdrix

OpenStudy (mathstudent55):

First work the numerator and denominator with a^m*a^n = a^(m + n) Then take care of the numerator divided by the denominator with a^m/a^n = a^(m - n) Then finally use the power raised to a power rule: (a^m)^n = a^(mn)